# Generative AI and Firm Productivity:

### v2

@@ 3.4 Data and Estimation
We obtained comprehensive granular consumer/product-level data for the first five of the seven experiments, allowing for in-depth measurement and exploration of productivity gains. For the remaining two experiments—Chargeback Defense and Live Chat Translation—the platform could not provide granular data. In these cases, we rely on analyses conducted by the platform’s internal data science team. These estimates complement our direct observations, offering a broader perspective on the impact of GenAI across various business areas (see “Data Availability” in Table 2).
For the experiments operated at the consumer level, we recorded each consumer’s treatment status. We gathered consumers’ complete set of activities, including the number of product views (referred to as View), product clicks (Click), product orders (Order), and total expenditures on those orders (Sales). [^13] In the Search Query Refinement experiment, product views represent the number of products a consumer browses on the search results page, which displays a summarized collection of products immediately after a query search. In the Google Advertising Title experiment, product views refer to the number of views of advertised products within the Google Shopping tab. In the Search Query Refinement and Product Descriptions experiments, product clicks capture the number of times consumers clicked into product detail pages. In the Marketing Push Message experiment, product clicks reflect consumer clicks on push notifications, while in the Google Advertising Title experiment, they indicate clicks on advertised products. For comparability across workflows, our benchmark analysis focuses on two primary outcome measures: (i) total consumer expenditure (Sales), which serves as a revenue-based measure of retail productivity; and (ii) the conversion rate, a binary indicator for whether a consumer made a purchase during the experimental period, representing a proxy for changes in consumer experience. Both sales and conversion rate are widely used industry metrics. For the product-level experiment, we collected analogous data at the product level. Table 3 presents summary statistics for the key variables in each of the five experiments where granular data are available.
To compare the mean purchase value and mean conversion rate between the treatment and control groups, we rely on the following general empirical specification, which we adapt as needed for each experiment:
@@
Conversion Rate 0.068 0.253 0 0 1
Sales 1.86 9.749 0 0 517.34
@@
View 313.36 615.02 125 1 105,883
Click 8.23 16.99 3 0 2,024
Order 0.16 0.73 0 0 85
@@
Click 0.017 0.131 0 0 1
Order 0.0018 0.053 0 0 6
Conversion Rate 0.0016 0.040 0 0 1
Sales 0.025 1.510 0 0 501
@@
View 19.36 82.38 5 2 12,033
Click 0.22 1.69 0 0 627
Conversion Rate 0.004 0.069 0 0 6
Sales 0.129 2.97 0 0 322
@@
    “View” refers to the number of product views. “Click” stands for the number of product clicks. “Order” is the number of product orders. “Sales” represents the total expenditure on product orders. “Conversion rate” measures consumers’ likelihood of making purchases. It is a binary indicator for purchase, which equals 1 if a consumer makes at least one order during the experiment period, and 0 otherwise. In Google Advertising Title experiment, the unit of observation is at the product level. The conversion rate is calculated as the number of purchases divided by the number of views for that product.
We estimate Equation (2) via OLS, adjusting the standard errors for heteroskedasticity. Under random assignment, $\beta$ recovers the average treatment effect of GenAI adoption, expressed as the absolute lift in outcomes. We also report results in percent lift, rescaling $\beta$ by the control group mean. For sales, we use levels to address concerns regarding log transformations with zero outcomes [19]. For conversions (a binary outcome), we primarily estimate a linear probability model and confirm that the findings are robust to logit specifications. We also estimate the model using pre-experiment covariates as controls, and the results remain consistent. Consumer overlap across experiments was minimal (less than 1%), and our findings are robust to excluding overlapping observations. This design allows treatment effects to be solely attributed to individual workflows, though it does not capture potential complementarities across GenAI applications.
@@
To enrich our analysis, we obtained pre-experiment seller data for products included in the experiments (e.g. seller size measured by annual sales, seller operational years, and the number of sub-accounts linked to a seller’s online store). These data enable us to examine seller-level heterogeneity in the three experiments involving products sold by both third-party sellers and the platform (see column “Product Sold By” in Table 2). However, this analysis is not applicable to the Pre-sale Service Chatbot and Product Description experiments, as the products in these cases were platform self-sold products—sold exclusively by a limited number of platform-operated sellers—resulting in insufficient variation in seller characteristics for meaningful heterogeneity analysis. Additionally, we collect product characteristics, such as the concentration ratio of its associated category, price, and annual sales quantity, to explore product-level heterogeneity.
We also augmented the data with consumer demographics and pre-experiment shopping history (years of registration, activity level, purchase volume, etc.), which supported both the analysis of buyer-level heterogeneity and the verification of random group assignment in the experiments. [^14] According to our research agreement with the partner platform, all consumers in our data are anonymous to ensure consumer privacy. We identify consumers by hashed IDs instead of knowing their actual names. As confirmed by the covariate balance checks reported in Figure 1, we found no systematic significant pre-experiment differences across consumers between the control and treatment groups. Thus, the randomization process was effective at allocating comparable consumers/products into the two groups. Details on the covariate balance checks are presented in Appendix B.
@@ 4 Main Results
Table 4 reports average effects across the GenAI-driven business workflows. The table presents, from left to right: (i) the impact of GenAI re-engineering on sales, our measure of productivity; and (ii) the mechanism underlying productivity gains, captured by the conversion rate as a proxy for consumer experience. Columns (1) and (3) report the estimated average treatment effects (ATE, absolute lift), while columns (2) and (4) present percentage changes relative to the control group (relative lift). Additional results are provided in Appendix C.
@@ 4.1 Productivity Impact by Workflow
Table 4 shows overall productivity improvements from most GenAI deployments, alongside substantial heterogeneity across workflows. The largest effect arises in the Pre-sale Service Chatbot, where GenAI increases sales by 16.3% ($p<0.01$) relative to the control group (Column 2). In this setting, consumers in the treatment group were assisted by a GenAI-powered chatbot, while those in the control group received an automated message indicating that no support was available. This auto-response condition reflects the platform’s standard practice of allocating human agents primarily to urgent post-sale inquiries, while offering no assistance for most pre-sale inquiries on self-sold products due to limited resources. However, it is plausible that consumers in the control group became frustrated by the lack of assistance, potentially reducing their likelihood of purchase and leading to an overestimation of our treatment effects. For this reason, the platform conducted further experiments on the GenAI-powered chatbot.
Appendix Table C1 reports additional comparisons, including evaluations against and in combination with human agents. The results show that the GenAI chatbot delivers service quality comparable to human customer support (Columns 3–4). Integrating GenAI with human agents further increases sales: when comparing the no pre-sale service condition with the treatment that combines GenAI assistance and human escalation when needed, sales rise by 25% (Columns 5–6), indicating strong complementarities between GenAI and human labor. More importantly, comparing consumers who receive GenAI-assisted service with human escalation to those served exclusively by human agents shows that the former spend 11.5% more (Columns 7–8). We interpret this latter comparison as a conservative lower bound on GenAI’s productivity impact in pre-sale customer support.
Table 4: Summary of Average Treatment Effects of GenAI Adoption Across Workflows
Productivity Impact (Sales, \$) Mechanisms (Conversion Rate)
@@
Business Workflow Coefficient % Change Coefficient % Change
1
@@
6 Chargeback Defense 15% defense success rate increase
7 Live Chat Translation 5.2% consumer satisfaction increase
@@ 6 Chargeback Defense 15% defense success rate increase
    “Sales” represents the total expenditure on product orders, in USD. “Conversion Rate” measures consumers’ likelihood of making purchases. It is a binary indicator for purchase, which equals 1 if a consumer makes at least one order during the experiment period, and 0 otherwise.
@@
    Columns (1) and (3) report the estimated coefficients for sales and conversion rate, respectively, with standard errors in brackets. Columns (2) and (4) report % Change for sales and conversion rate, respectively. % Change is calculated by dividing the treatment effect by the control group average. *** p<<0.01, ** p<<0.05, * p<<0.1.
  • 3
The remaining four workflows with detailed data reported in Table 4 exhibit smaller effects on sales, ranging from a negative and statistically insignificant change to gains of up to 3%. Specifically, the Search Query Refinement application increased sales by 2.93% ($p<0.05$), while the Product Description generated a 2.05% ($p<0.05$) gain—still substantial effects for a platform of this scale and maturity. The Marketing Push Message workflow shows a positive yet not statistically significant improvement in sales (1.6%). This likely reflects the combination of a very low baseline conversion rate (only 0.14% of consumers make a purchase) and high variance in expenditures among converters, suggesting that broader implementation could provide sufficient power to detect treatment effects. [^15] Note that this case exhibits significant conversion rate increase (3.1%) in Table 4 as well as significant order increase (2.8%) in Appendix Table C5. It is also important to note that only a subset of the treatment group was exposed to GenAI-generated messages in this experiment, which may attenuate the observed treatment effects under partial exposure. By contrast, the Google Advertising Title workflow exhibits an insignificant negative effect. This pattern is consistent with the lack of fine-tuning of the GenAI model to the advertising context within the e-commerce domain, which led it to omit commonly used commercial keywords. An alternative explanation is a divergent delivery bias in ad distribution: Google’s advertising algorithm may have recognized and deprioritized AI-generated titles, reducing their visibility. Together, these findings underscore the importance of domain-specific fine-tuning or retraining of foundation models when applying GenAI to industry-specific tasks requiring specialized knowledge [27].
For the last two business processes studied (Chargeback Defense and Live Chat Translation), we did not obtain granular transaction data and instead relied on the company’s internal metrics and analysis. While not directly comparable to the effects on sales, these measures indicate substantial improvements in two important outcomes: a 15% increase in defense success rates for Chargeback Defense and a 5.2% increase in consumer satisfaction attributable to Live Chat Translation. [^16] In Chargeback Defense, further gains not captured in our calculations may also arise from cost reductions, as the GenAI-enhanced workflow eliminated the need for manual intervention.
These results provide new evidence on the potential of GenAI to enhance revenue-based productivity in online retail, thereby contributing to the broader debate on the economic consequences of GenAI adoption [56, 1]. A key distinction of our study is its focus on revenue outcomes rather than input-side efficiency gains, which makes our estimates not directly comparable to those from studies emphasizing labor productivity or input-side task efficiency. Among the seven deployments we examine, most deliver measurable performance gains, showing that GenAI can generate substantial improvements in firm outcomes under real-world operating conditions. Overall, the evidence points to a positive but heterogeneous impact of GenAI along the customer journey.
The variation we observe across workflows is unlikely to reflect differences in implementation quality—which was comparable across applications—but instead arises from differences in the scope of GenAI’s marginal contribution relative to baseline conditions. In each case, the control group reflects the firm’s standard practices prior to GenAI adoption. For example, in the Pre-Sale Service Chatbot, the control group received no customer support; in Search Query Refinement, it relied on standard machine-learning–based search algorithms; and in most other workflows, the benchmark was human input. The results therefore point to genuine heterogeneity in where GenAI is most effective: customer-support applications, such as Pre-Sale Service Chatbots, deliver the largest improvements; search and product-discovery tasks yield smaller gains; and advertising-related applications show no significant effects. [^17] In the Marketing Push Message workflow, experimental design features may also influence measured effects, as only a subset of the treated population was exposed to GenAI, which likely attenuated measured impacts. Together, this highlights both the role of baseline conditions in shaping treatment effects and the comparative effectiveness of GenAI across functional areas.
4.2 Mechanism
Most of the existing literature attributes the productivity-enhancing potential of GenAI to supply-side mechanisms, such as labor savings or efficiency improvements (e.g., reductions in the time required to complete a task). Our context and findings highlight an additional channel: GenAI can unlock productivity gains by enriching the demand-side consumer experience. We measure this mechanism using conversion rates (i.e., the likelihood that consumers complete a purchase at the extensive margin), which serve as a widely recognized industry standard for consumer satisfaction in the e-commerce context. Across the various workflows, we document significant increases in conversion rates ranging from 1% to 22% (Column 4 of Table 4), which in turn translate into higher output as measured by sales. Notably, we do not find any significant effects along the intensive margin. Table 5 shows that the average cart value among consumers who made at least one purchase (or among products purchased at least once) remains unchanged following GenAI adoption. In our setting, GenAI primarily drove market expansion by converting a larger share of consumers, rather than inducing existing buyers to purchase higher-priced or larger quantities of products.
Table 5: Impact of GenAI Adoption on Average Cart Value (intensive margins)
Cart Value (\$)
(1) (2)
Business Workflow Coefficient % Change
1 Pre-sale Service Chatbot -0.859 -3.1%
2 Search Query Refinement 0.376 1.49%
3
Product Description 0.0942 0.81%
4
Marketing Push Message 0.024 0.15%
5
Google Advertising Title -0.784 -2.3%
@@ 4.2 Mechanism
    The table presents the treatment effects of GenAI adoption on “Cart Value”. For workflows 1-4, “Cart Value” refers to the expenditure per consumer, conditional on the consumer making a purchase. For workflow 5, “Cart Value” refers to the sales per product, conditional on the product being purchased.
@@
    Column (1) reports the estimated coefficient, with standard errors in brackets. Column (2) reports % Change, which is calculated dividing the treatment effect by the control group mean. *** p<<0.01, ** p<<0.05, * p<<0.1.
This evidence indicates that the observed productivity gains arise from GenAI’s capacity to reduce market frictions—both by enabling new services and by improving existing ones—thereby enhancing the shopping experience and influencing purchasing behavior. For example, GenAI mitigates information asymmetries by providing relevant and timely assistance through pre-sale chatbots (a 21.7% increase in conversion rate, $p<0.01$) and by generating comprehensive and structured product descriptions (a 1.27% increase, $p<0.01$). It reduces search frictions and improves match quality by enhancing the translation and semantic comprehension of consumer queries (a 1.15% increase, $p<0.05$). In addition, GenAI enables personalization of marketing content by enabling large-scale generation of customized marketing messages across a broad product portfolio (a 3% increase, $p<0.05$). Finally, evidence from Chargeback Defense and Live Chat Translation indicates improvements consistent with enhanced experiences for sellers and consumers, respectively, including higher dispute success rates and greater customer satisfaction.
In Appendix C, we provide additional results on user activities where data are available. For example, GenAI-refined search queries led to a 2.02% $(p<0.01)$ increase in click-through rates—the ratio of product clicks to views—indicating that consumers found the displayed products more appealing and were more likely to seek additional information after viewing the summarized search results. GenAI-generated product descriptions increased the number of orders placed by 1.08% $(p<0.05)$, while GenAI-created marketing messages raised clicks by 3% $(p<0.05)$ and orders by 2.8% $(p<0.10)$. These additional outcomes provide consistent evidence that GenAI enhances intermediate engagement measures—such as clicks, orders, and click-through rates—reflecting a smoother and more informative shopping experience that, in turn, underpins the higher conversion and sales gains documented above.

### v3

@@ 3.4 Data and Estimation
We obtained comprehensive granular consumer- or product-level data for the first five of the seven experiments, allowing for in-depth analysis of productivity gains. For the remaining two experiments—Chargeback Defense and Live Chat Translation—the platform could not provide granular data. In these cases, we rely on analyses conducted by the platform’s internal data science team. These estimates complement our direct observations, offering a broader perspective on the impact of GenAI across various business areas (see “Data Availability” in Table 2).
For experiments conducted at the consumer level, we record each consumer’s treatment status and observe their complete set of activities, including the number of product views (Views), product clicks (Clicks), product orders (Orders), and total expenditure on those orders (Sales). [^14] In Search Query Refinement, product views represent the number of products a consumer browses on the search results page, which displays a summarized collection of products immediately after a query search. In Google Advertising Title, product views refer to the number of views of advertised products within the Google Shopping tab. In Search Query Refinement and Product Descriptions, product clicks capture the number of times consumers clicked into product detail pages. In Marketing Push Message, product clicks reflect consumer clicks on push notifications, while in Google Advertising Title, they indicate clicks on advertised products. For comparability across workflows, our main analysis focuses on sales, measured as total consumer expenditure, which serves as our primary revenue-based measure of productivity. To shed light on the mechanisms underlying the sales effects, we additionally examine outcomes along the extensive and intensive margins. In particular, we analyze conversion rates, defined as a binary indicator for whether a consumer makes at least one purchase during the experimental period, and cart value, measured as average expenditure conditional on making a purchase. Conversion rates capture changes in purchase incidence and serve as a widely used proxy for consumer experience in online retail, while cart value reflects adjustments along the intensive margin. For the product-level experiment, we collect the same outcomes at the product level. Table 3 reports summary statistics for these key variables across the five experiments for which granular data are available.
To compare mean sales, conversion rates, and cart values between the treatment and control groups, we use the following general empirical specification, adapted as needed for each experiment:
@@
We estimate Equation (2) via OLS, adjusting the standard errors for heteroskedasticity. Under random assignment, $\beta$ recovers the average treatment effect of GenAI adoption, expressed as the absolute lift in outcomes. We also report results in percent lift, rescaling $\beta$ by the control group mean. For sales, we use levels to address concerns regarding log transformations with zero outcomes [undefr]. For conversions (a binary outcome), we primarily estimate a linear probability model and confirm that the findings are robust to logit specifications. We also estimate the model using pre-experiment covariates as controls, and the results remain consistent. Consumer overlap across experiments was minimal (less than 1%), and our findings are robust to excluding overlapping observations. This design allows treatment effects to be solely attributed to individual workflows, though it does not capture potential complementarities across GenAI applications.
@@
Sales 1.86 9.75 0 0 517
Conversion Rate 0.07 0.25 0 0 1
Cart Value 26.71 26.21 18.61 1.11 517
@@
Views 313.36 615.02 125 1 105,883
Clicks 8.23 16.99 3 0 2,024
Orders 0.16 0.73 0 0 85
Sales 2.24 21.41 0 0 4,960
@@
Clicks 0.017 0.13 0 0 1
Orders 0.0018 0.05 0 0 6
Sales 0.025 1.51 0 0 501
Conversion Rate 0.0016 0.04 0 0 1
Cart Value 15.52 33.97 7.81 0.01 501
@@
Views 19.36 82.38 5 2 12,033
Clicks 0.22 1.69 0 0 627
Sales 0.13 2.97 0 0 322
Conversion Rate 0.004 0.07 0 0 6
Cart Value 41.35 72.16 22.06 0.54 322
@@
    “Views” refers to the number of product views. “Clicks” denotes the number of product clicks. “Orders” is the number of product orders. “Sales” represents the total expenditure on product orders. “Conversion rate” measures consumers’ likelihood of making a purchase. It is a binary indicator for purchase, which equals 1 if a consumer makes at least one order during the experiment period, and 0 otherwise. “Cart value” is the expenditure per consumer, conditional on the consumer making a purchase.
  • 1
    In the Google Advertising Title experiment, the unit of observation is the product. The conversion rate is calculated as the number of purchases divided by the number of views for a given product, while cart value refers to sales per product, conditional on the product being purchased.
To enrich our analysis, we obtained pre-experiment seller data for products included in the experiments (e.g. seller size measured by annual sales, operational years, and the number of sub-accounts linked to a seller’s online store). These data enable us to examine seller-level heterogeneity in the three experiments involving products sold by both third-party sellers and the platform (see column “Product Sold By” in Table 2). This analysis is not applicable to the Pre-sale Service Chatbot and Product Description experiments, as both involve platform self-sold products sold by a limited number of platform-operated sellers, resulting in insufficient variation in seller characteristics for meaningful heterogeneity analysis. Additionally, we collect product characteristics, such as the concentration level of the associated category, price, and annual sales quantity, to explore product-level heterogeneity.
We also augmented the data with consumer demographics and pre-experiment shopping history (years of registration, activity level, purchase volume, etc.), which supported both the analysis of buyer-level heterogeneity and the verification of random group assignment in the experiments. [^15] According to our research agreement with the partner platform, all consumers in our data are anonymous to ensure consumer privacy. We identify consumers by hashed IDs instead of knowing their actual names. As confirmed by the covariate balance checks reported in Figure 1, we found no systematic significant pre-experiment differences across consumers between the control and treatment groups. Thus, the randomization process was effective at allocating comparable consumers and products into the two groups. Details on the covariate balance checks are presented in Appendix B.
@@ 4.1 Productivity Impact by Workflow
We begin by examining the impact of GenAI on productivity across business workflows. Table 4 reports the estimated effects on sales, our primary productivity measure. Column (1) presents the average treatment effect (ATE), capturing the absolute change in sales, while column (2) reports the corresponding percentage change relative to the control group. The results presented here do not account for differences in the frequency of consumer interactions with the workflows or in the duration of the corresponding experiments. For example, search affects nearly all consumers on a daily basis, whereas chatbot interactions occur less frequently. In Section 4.3, we provide adjusted measures that account for these differences. Additional results and robustness analyses are reported in Appendix C.
Table 4: Average Treatment Effects of GenAI Adoption on Sales Across Workflows
Productivity Impact (Sales, \$)
@@
Business Workflow Coefficient % Change Obs Unit of Obs
@@
1,244,016
Product
Chargeback Defense 15% defense success rate increase
Live Chat Translation 5.2% consumer satisfaction increase
  • 1
    “Sales” represents the total expenditure on product orders, in USD.
  • 2
    Columns (1) reports the estimated coefficients, with standard errors in parentheses. Columns (2) reports % Change, calculated as the treatment effect divided by the control group mean. Columns (3) reports the number of observations. *** p<<0.01, ** p<<0.05, * p<<0.1.
  • 3
The table shows that most GenAI deployments are associated with productivity gains, with substantial heterogeneity across workflows. The largest effect is observed in the pre-sale service chatbot workflow, where GenAI increases sales by 16.3% relative to the control group ($p<0.01$; Column 2). In this setting, consumers in the treatment group interact with a GenAI-powered chatbot, whereas consumers in the control group receive an automated message indicating that real-time support is unavailable. This control condition reflects the platform’s standard operating practice: human agents are prioritized for post-sale inquiries, leaving a substantial share of pre-sale inquiries for self-sold products without live assistance due to capacity constraints. This setting also mirrors the operating reality faced by many third-party sellers on the platform, particularly smaller-scale sellers who lack the capacity to provide real-time, multilingual customer support. Taken together, these features suggest that GenAI enhances productivity by performing tasks that are difficult for human agents to fulfill due to talent shortages or resource constraints.
One potential concern is that consumers in the control group may have been frustrated by the absence of assistance, mechanically reducing their likelihood of purchase and thus inflating the estimated treatment effects. To address this concern, we conducted additional experiments on the GenAI-powered chatbot. Appendix Table C1 reports a series of comparisons, including benchmarks against human agents and hybrid configurations that combine GenAI assistance with human escalation when needed. The results indicate that the GenAI chatbot delivers service quality comparable to human customer support (Column 2). Moreover, integrating GenAI with human agents yields substantially larger effects: relative to the no pre-sale service condition, sales increase by 25% when GenAI assistance is combined with human escalation ($p<0.01$; Columns 3), pointing to strong complementarities between GenAI and human labor. Most importantly, comparing consumers who receive GenAI-assisted service with human escalation to those served exclusively by human agents shows that the former spend 11.5% more ($p<0.1$; Columns 4). We interpret this comparison as a conservative lower bound on GenAI’s productivity impact in pre-sale customer support.
The remaining four workflows with detailed data reported in Table 4 exhibit more modest effects on sales, ranging from a negative and statistically insignificant estimate to gains of up to 3%. In particular, the Search Query Refinement application increases sales by 2.9% ($p<0.05$), while automated Product Description generation raises sales by 2.1% ($p<0.05$). Although smaller in magnitude, these effects are economically meaningful for a platform of this scale and maturity, especially given the near-universal exposure of consumers to search and product descriptions.
The Marketing Push Message workflow shows a positive yet not statistically significant improvement in sales (1.6%). This likely reflects the combination of a very low baseline conversion rate (only 0.16% of consumers make a purchase) and high variance in expenditures among converters, suggesting that broader implementation could provide sufficient power to detect treatment effects. Notably, however, this case exhibits significant increases in conversion rate and number of orders, which we discuss in detail later. It is also important to note that, for risk-mitigation purposes, only a subset of the treatment group was exposed to AI-generated messages in this experiment, which may attenuate the observed treatment effects under partial exposure.
By contrast, the Google Advertising Title workflow exhibits a statistically insignificant negative effect. This pattern is consistent with the GenAI model not being fine-tuned to the advertising context within e-commerce, leading it to omit commercially salient keywords commonly used in ad titles. As a result, these ads may receive lower quality scores by Google, reducing their probability of winning auctions at a given bid. In addition, Google’s advertising algorithm may identify and deprioritize AI-generated titles, further lowering their visibility. Consistent with this interpretation, Table C8 in the Appendix shows that ads with AI-generated titles receive fewer views and clicks than those with human-generated titles. Taken together, these results underscore the importance of domain-specific fine-tuning or retraining of foundation models when deploying GenAI in industry-specific tasks that require specialized knowledge [undefz].
For the final two business processes studied—Chargeback Defense and Live Chat Translation—we do not observe granular transaction-level data and instead rely on the platform’s internal metrics and analyses, which do not report statistical significance. While these outcomes are not directly comparable to the sales effects discussed above, they nevertheless point to substantial improvements in two operational dimensions. Specifically, Chargeback Defense is associated with a 15% increase in defense success rates, and Live Chat Translation increases consumer satisfaction by 5.2%. Both effects point directionally to successful implementation of the GenAI applications. [^16] In the case of Chargeback Defense, additional gains not captured in our calculations may arise from cost reductions, as the GenAI-enabled workflow eliminates the need for manual intervention.
These results provide new evidence on the potential of GenAI to enhance revenue-based productivity in online retail, thereby contributing to the broader debate on the economic consequences of GenAI adoption [undefaac, undef]. A key distinction of our study is its focus on revenue outcomes rather than input-side efficiency gains, which makes our estimates not directly comparable to those from studies emphasizing labor productivity or input-side task efficiency. Among the seven deployments we examine, most deliver measurable performance gains, showing that GenAI can generate substantial improvements in firm outcomes under real-world operating conditions. Overall, the evidence points to a positive but heterogeneous impact of GenAI along the customer journey.
The variation in effects across workflows is unlikely to reflect differences in implementation quality, which was comparable across applications. Instead, it appears to arise from differences in the marginal contribution of GenAI relative to baseline conditions. In each case, the control group reflects the firm’s standard operating practices prior to GenAI adoption. For example, in the Pre-Sale Service Chatbot workflow, the control group received no customer support; in Search Query Refinement, it relied on standard machine-learning-based search algorithms with lexical translation into English; and in most other workflows, the benchmark was human input. The results therefore point to genuine heterogeneity in where GenAI is most effective. Customer-support applications, such as Pre-Sale Service Chatbots, generate the largest improvements; search and product-discovery tasks yield more modest gains; and advertising-related applications exhibit no statistically significant effects. [^17] In the Marketing Push Message workflow, features of the experimental design may also affect measured impacts, as only a subset of treated users was exposed to GenAI, likely attenuating estimated effects. Taken together, these findings highlight both the role of baseline conditions in shaping treatment effects and the differential effectiveness of GenAI across functional areas.
4.2 Mechanisms
Much of the existing literature attributes the productivity-enhancing potential of GenAI to input-side mechanisms, such as labor savings or efficiency improvements that reduce the time required to complete tasks. In this section, we examine an additional channel operating on the demand side: GenAI can increase productivity by improving the consumer experience during the shopping process.
We investigate two primary mechanisms behind the observed impact on sales. The first is market expansion, captured by an increase in the probability that a consumer completes a purchase at the extensive margin. We measure this mechanism using conversion rates, a widely used industry metric that reflects consumers’ revealed satisfaction and perceived friction during the ex ante search and evaluation process in e-commerce. The second mechanism is a change in the behavior or preferences of consumers who make a purchase, reflected in adjustments to spending intensity and order composition. In this case, GenAI may influence consumers’ willingness to pay, alter purchasing patterns, or shift demand toward products at different price points. We assess this channel using cart value and purchase frequency, conditional on consumers making a purchase.
Table 5: Average Treatment Effects of GenAI Adoption on Conversion Rate and Cart Value
Conversion Rate Cart Value (\$)
(1) (2) (3) (4) (5) (6)
Business Workflow Coefficient % Change Obs Coefficient % Change Obs
Pre-sale Service Chatbot
0.0131***
21.7%
44,614
-1.264
-4.5%
3,076
Search Query Refinement
0.00101**
1.1%
1,849,382
0.370
1.5%
163,381
Product Description
0.000554***
1.3%
4,772,937
0.0944
0.8%
210,155
Marketing Push Message
0.000048**
3.0%
13,715,528
-0.206
-1.3%
22,425
Google Advertising Title
@@ 4.2 Mechanisms
1,244,016
-0.784
-2.3%
4,811
@@
    “Conversion Rate” measures consumers’ likelihood of making purchases. It is a binary indicator for purchase, which equals 1 if a consumer makes at least one order during the experiment period, and 0 otherwise. For workflows 1-4, “Cart Value” refers to the expenditure per consumer, conditional on the consumer making a purchase. For workflow 5, “Cart Value” refers to the sales per product, conditional on the product being purchased.
@@
    Columns (1) and (4) report the estimated coefficients, with standard errors in parentheses. Columns (2) and (5) report % Change, calculated as the treatment effect divided by the control group mean. Columns (3) and (6) report the number of observations. *** p<<0.01, ** p<<0.05, * p<<0.1.
Table 5 reports the results on conversion rates and cart value, with additional evidence on purchase frequency provided in Appendix C. Across workflows, we document economically and statistically significant increases in conversion rates ranging from 1% to 22% (Column 2), which translate directly into higher output as measured by sales. By contrast, we find no evidence of effects along the intensive margin. Columns 4 and 5 show that cart value, as measured by the average order value among consumers who make at least one purchase (or among products purchased at least once, in the case of Google Advertising Title), remains unchanged following GenAI adoption. We likewise find no significant changes in purchase frequency, defined as the number of orders placed conditional on making at least one purchase (Appendix C). Taken together, these results indicate that GenAI primarily drives productivity gains through market expansion rather than by altering spending patterns among existing buyers.
The evidence on increased conversion rates in Table 5 suggests that GenAI has the potential to reduce market frictions, both by enabling new services and by improving existing ones, thereby enhancing the shopping experience and increasing the probability of purchase. For example, GenAI can mitigate search costs and information asymmetries by providing relevant and timely assistance through a pre-sale chatbot (a 21.7% increase in conversion rates, $p<0.01$) and by generating more comprehensive and structured product descriptions (a 1.3% increase, $p<0.01$). It can also reduce search frictions and improve match quality by enhancing the translation and semantic understanding of consumer queries (a 1.1% increase, $p<0.05$). In addition, GenAI enables personalization of marketing content at scale by generating customized messages across a broad product portfolio (a 3% increase, $p<0.05$). Finally, evidence from Chargeback Defense and Live Chat Translation reported in Table 4 points to improvements consistent with enhanced experiences for sellers and consumers, respectively, including higher dispute success rates and greater customer satisfaction.
In Appendix C, we report additional evidence on how GenAI reduces frictions along the customer journey, where such data are available. For example, Table C4 shows that GenAI-refined search queries increase the likelihood of a product click by 0.3% ($p<0.01$; Column 1), consistent with improved search performance in combination with the previously documented increase in purchase conversion. Columns 2 and 3 further indicate that treated consumers view fewer products prior to clicking or purchasing, suggesting a reduction in search intensity. These patterns align with prior literature showing that improvements in search quality increase conversion while reducing search effort [undefaal, undefaap]. In addition, we observe a statistically significant 2.0% increase in the click-through rate ($p<0.01$; Column 4), defined as the ratio of product clicks to product views, while the click-to-order conversion rate (Column 5), defined as the ratio of orders to clicks, remains statistically insignificant. Together, these findings suggest that GenAI-refined queries primarily affect the composition of products retrieved at the search stage, rather than the information displayed on product detail pages, reinforcing the proposed query-refinement mechanism.
Table C5 shows that AI-generated product descriptions increase the number of orders placed by 1.1% ($p<0.05$; Column 2). Moreover, when we stratify products by the length of their original, human-generated descriptions, we find substantial heterogeneity in effects. Products with no or insufficient textual descriptions (fewer than 50 words) experience a 6.1% increase in sales following augmentation with AI-generated content ($p<0.05$), whereas products with sufficiently detailed descriptions (more than 50 words) exhibit no significant increase (Columns 1 and 2 of Table C6). [^18] Based on internal research and expert surveys conducted by our partner platform, descriptions containing fewer than 50 words are classified as providing insufficient or minimal textual information. This pattern indicates that the descriptions augmented by GenAI are most effective for products with limited baseline information.
In the marketing push workflow, GenAI enables large-scale variation in customer-facing content. The treatment group includes approximately 2.7 million unique message variants, compared with only roughly 2,000 variants in the control group, which relies exclusively on human-generated content. Consistent with the expanded scope for personalization, AI-generated marketing messages increase clicks by 3.1% ($p<0.05$) and orders by 2.8% ($p<0.10$; Columns 2 and 3 of Table C7). Taken together, these results provide consistent evidence that GenAI enhances intermediate engagement outcomes, such as clicks, orders, and click-through rates, reflecting a smoother and more informative shopping experience that underpins the higher conversion and sales gains documented above.
A natural concern with deploying GenAI in customer-facing processes is that it may introduce hallucinations or distortions that mislead consumers—for example, by inducing impulse purchases or manipulation—thereby reducing ex post consumer satisfaction. In our case, increases in revealed purchase behavior may not map cleanly into welfare gains, as higher sales could reflect distorted choices rather than improved consumer surplus. To assess this possibility, we examine two post-purchase outcomes that capture consumers’ evaluations of their purchase decisions. The first is the product return rate, defined as the share of orders that are returned. The second is the positive review rate, defined as the share of rated orders that receive four- or five-star ratings on a five-point scale.
Figure 2: Average Treatment Effects of GenAI Adoption on Product Returns and Reviews
@@
    “Return Rate” is defined as the share of orders that are returned. “Positive Review Rate” is defined as the share of rated orders that receive four- or five-star ratings on a five-point scale. The estimation of Return Rate is conditional on consumers with orders, while the estimation of Positive Review Rate is conditional on consumers provide ratings. Such data are not available for the Google Advertising Title workflow.
@@
    Full regression results are reported in Table C9. *** p<<0.01, ** p<<0.05, * p<<0.1.
As shown in Figure 2 (full regression results are provided in Table C9), we find no evidence that the observed sales gains come at the expense of ex post consumer satisfaction. Product return rates and customer ratings do not deteriorate following GenAI adoption. Notably, return rates decline significantly for Product Descriptions and Marketing Push Messages (by 3.9% and 11.4%, respectively; $p<0.01$), consistent with improvements in both ex ante decision quality and ex post consumer satisfaction in these workflows. We also observe a positive effect on customer ratings for the Pre-sale Service Chatbot (a 4.5% increase; $p<0.10$). In all other cases, the estimated effects are statistically insignificant, providing no evidence that GenAI induces deceptive practices that mislead consumer decisions.